OpenStudy (anonymous):

Two metals X and Y are to be used for making two different alloys. If the ratio by weight of X : Y in the first alloy is 6 : 5 and that in the second is 7 : 13, how many kg of X metal must be melted along with 11 kg of the first alloy and 20 kg of the second so as to produce a new alloy containing 40% of metal Y?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

6x+5y=11 7x+13y=20

OpenStudy (anonymous):

x=1 y=1

OpenStudy (anonymous):

x=13 y=18

OpenStudy (anonymous):

18*40/100=36/5

OpenStudy (anonymous):